OpenStudy (anonymous):

Anti derivative of sqrt(t^11)

OpenStudy (johnweldon1993):

\[\large \int\limits_{}^{}\sqrt{t^{11}}\] like that right?

OpenStudy (anonymous):

yeah

OpenStudy (johnweldon1993):

Sorry lets keep it in terms of 't' Alright...so do a u-sub...let u = √t this would make du = 1/(2√t)dt can you take it from here?

OpenStudy (phi):

you could rewrite it as \( t^\frac{11}{2}\)

OpenStudy (anonymous):

then can I do t^n+1/ n+1

OpenStudy (phi):

yes